- 170: KH2O, Site All, Tue 19-Dec-2006 19:41:27 MST, Krypton is Quality Controlled
All the kryptons have been quality controlled.
This includes removal of rain dropouts, cleanings, malfunctioning sensor,
and sensor replacement.
- 141: KH2O, Site Southwest, Thu 26-Oct-2006 17:38:23 MST, generator-spike
There is a -Inf spike on h2o at about 1100 today. This is when the generator
was being serviced. We thought that the battery at the station would keep
it going through this, but apparently the station voltage dropped to 11.4V
during this period.
It looks like it came back up okay.
- 132: KH2O, Site East_Upper, Wed 25-Oct-2006 14:07:31 MST, krypton klean
krypton kleaned ~1248. MSB of counts went from 0d -> 42!
- 131: KH2O, Site East_Lower, Wed 25-Oct-2006 14:07:08 MST, krypton klean
krypton kleaned ~1217. MSB of counts went from 0e -> 1f.
- 130: KH2O, Site West_Upper, Wed 25-Oct-2006 14:06:42 MST, krypton klean
krypton kleaned ~1155. MSB of counts went from 09 -> 26.
Krypton mount was loose, but I didn't have allen wrenches. Should be revisited.
- 129: KH2O, Site West_Lower, Wed 25-Oct-2006 14:06:08 MST, krypton klean
krypton kleaned ~1145. MSB of counts went from 03 -> 23.
I had to restart data system code (adn/aup) to get rserial to work. (But it
seems like the archiving had been fine.)
- 128: KH2O, Site Floor, Wed 25-Oct-2006 14:04:51 MST, krypton klean
krypton kleaned ~1130. MSB of counts went from 08 -> 17.
- 50: KH2O, Site Floor, Mon 09-Oct-2006 14:29:37 MST, krypton changed at flr
krypton sensor s/n 1101 was replaced with s/n 1397
at approx. 1900 UTC 09 Oct. Unit 1101 was brought
back to the base.
- 32: KH2O, Site All, Fri 06-Oct-2006 19:24:25 MST, krypton serial numbers and calibration coefficients
Charlie entered these calibrations in xml configuration files, 10/7/06.